# On the Robustness of the Successive Projection Algorithm

### v1

@@ 1 Introduction
Countless algorithms have been developed for separable SSMF; see, e.g., [14, Chap. 7] and the references therein. Among them, a workhorse algorithm is the successive projection algorithm (SPA) [3] which is described in Algorithm 1. Geometrically, when $\operatorname{rk}(W)=r$, $\operatorname{conv}(W)$ is an $(r-1)$-dimensional simplex with $r$ vertices embedded in a higher $m$-dimensional space. The first step of SPA extracts the point with the largest $\ell_{2}$ norm. For any polytope (such as a simplex), such a point is always a vertex. Then SPA projects all data points on the orthogonal complement of that extracted point (so that this point is projected onto zero), reducing the dimension of the simplex by one, and continuing the process $(r-1)$ times.
@@
3:        $p\in\operatorname{argmax}_{j}\|R(:,j)\|_{2}$.    % Selection step
@@
5:        $R\leftarrow\left(I-\frac{{R(:,p)}R(:,p)^{\top}}{\|{R(:,p)}\|_{2}^{2}}\right)R$.      % Projection step
@@
SPA should not be implemented as in Algorithm 1, this would be particularly ineffective for large sparse matrices. It can be implemented in $\mathcal{O}(r\operatorname{nnz}(X))$ operations (essentially, $r$ matrix-vector products), where $\operatorname{nnz}(X)$ is the number of non-zero entries of $X$, using the fact that $\|(I-uu^{\top})y\|_{2}^{2}=\|y\|_{2}^{2}-(u^{\top}y)^{2}$ where $u$ has unit norm and $y$ is any vector [17]. See https://gitlab.com/ngillis/robustSPA for a MATLAB implementation.
SPA has a long history, as the index set of columns it extracts corresponds to that of QR with column pivoting [7]. However, it was introduced in the context of separable SSMF by Araújo et al. [3], and rediscovered many times under different names; see [14, p. 229] for an historical overview.
@@
[17, Theorem 3] Let $X=WH+N\in\mathbb{R}^{m\times n}$ where $H\in\mathbb{R}^{r\times n}_{+}$ is a separable matrix satisfying $H^{\top}e\leq e$, with $e$ the vector of all one of appropriate dimension, and let
$\varepsilon=\max_{j}\|N(:,j)\|_{2}\leq\mathcal{O}\left(\frac{\sigma_{r}(W)}{\sqrt{r}\mathcal{K}^{2}(W)}\right),$ (Bound on the noise)
where $\sigma_{r}(W)$ is the $r$th singular value of $W$, and $\mathcal{K}(W)=\frac{K(W)}{\sigma_{r}(W)}$ with $K(W)=\max_{k}\|W(:,k)\|_{2}$ is a measure of the conditioning [^2] Note that $\mathcal{K}(W)$ is smaller than the condition number of $W$, $\frac{\sigma_{1}(W)}{\sigma_{r}(W)}$. In fact, recall that $\sigma_{1}(W)=\|W\|_{2}\geq K(W)\geq\sigma_{1}(W)/\sqrt{r}$, where $r$ is the number of columns in $W$. of $W$. Then SPA (Algorithm 1) returns a set of indices $\mathcal{J}$ such that
$\max_{1\leq k\leq r}\min_{j}\|W(:,k)-X(:,\mathcal{J}_{j})\|_{2}\leq\mathcal{O}\left(\varepsilon\mathcal{K}^{2}(W)\right).$ (Error bound)
@@
  • Section 5. For $r\geq 3$, we show that the bound of Theorem 1 from [17] with error bound in $\mathcal{O}\left(\varepsilon\mathcal{K}^{2}(W)\right)$ is tight (Theorem 7). We do this by providing a family of data matrices that achieve this worst-case bound already at the second step of SPA. This answers an open question about the tightness of the SPA error bound of Theorem 1 [14, p. 231].
@@
Given a vector $x\in\mathbb{R}^{m}$, we denote $\|x\|$ or $\|x\|_{2}$ its $\ell_{2}$ norm. Given a matrix $X\in\mathbb{R}^{m\times n}$, we denote $X^{\top}$ its transpose, $X(:,j)$ its $j$th column, $X(i,:)$ its $i$th row, $X(i,j)$ its entry as position $(i,j)$, $\|X\|=\sigma_{\max}(X)$ its $\ell_{2}$ norm which is equal to its largest singular value, $\|X\|_{F}^{2}=\sum_{i,j}X(i,j)^{2}$ its squared Frobenius norm, $\sigma_{r}(X)$ its $r$th singular value, $\sigma_{\min}(X)=\sigma_{\min(m,n)}(X)$ its smallest singular value, $\operatorname{rk}(X)$ its rank, $K(X)$ $=$ $\max_{j}\{\|X(:,j)\|\}$, and $\mathcal{K}(X)=K(X)/\sigma_{\min}(X)$. For $k>\min\{m,n\}$, by convention, $\sigma_{k}(X)=0$. Note that
@@
and $\sigma_{n}(X)\sqrt{2}\leq\|X(:,i)-X(:,j)\|$ for every $i\neq j$, where $P_{u}(v)$ is the projection of $v$ on the subspace orthogonal to $u$.
We denote $e_{k}$ the $k$th unit vector, $e$ the vector of all ones of appropriate dimension, $I_{r}$ the identity matrix of dimension $r$. The set +$\mathbb{R}^{m\times n}_{+}$ denotes the $m$-by-$n$ component-wise nonnegative matrices. A matrix $H\in\mathbb{R}^{r\times n}$ has stochastic columns if $H\geq 0$ and $H^{\top}e=e$. It is separable if $H(:,\mathcal{K})=I_{r}$ for some index set $\mathcal{K}$ of size $r$.
@@ 2 SPA: what do we know?
    The error bound are rather poor, depending on the squared of the condition number of $W$. Moreover, it is not known whether this bound is tight [14, p. 231].
@@ 2.1 Drawback 1: rank-deficient case
3:        $p\in\operatorname{argmax}_{j}\|R(:,j)\|_{2}$.                 % Selection step
@@
8:              $R\leftarrow\left(I-\frac{{R(:,p)}R(:,p)^{\top}}{\|{R(:,p)}\|_{2}^{2}}\right)R$.      % Projection step
@@
T-SPA should not be implemented as in Algorithm 2, this would be particularly ineffective for large sparse matrices, as the residual would become quickly dense (especially if the first extracted columns is dense). As for SPA, it can be implemented in $\mathcal{O}(r\operatorname{nnz}(X))$ operations (essentially, $r$ matrix-vector products). See https://gitlab.com/ngillis/robustSPA for a MATLAB implementation.
@@
    Note that we will always have $\mathcal{K}(W_{t})\leq 2\mathcal{K}(W)$. In fact, $K(W_{t})\leq 2K(W)$ since $\|w_{i}-w_{j}\|_{2}\leq\|w_{i}\|_{2}+\|w_{j}\|_{2}$, and $\sigma_{{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}r}}-1}(W_{t})\geq\sigma_{{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}r}}}(W)$ since $W_{t}$ is a rank-one correction of $W$ (by the interlacing singular value theorem).
@@ 3.1 First step of SPA
[17, Theorem 2] Let $W\in\mathbb{R}^{m\times k}$, and $Q\in\mathbb{R}^{m\times(r-k)}$ and $X=[W\,\,Q]H+N\in\mathbb{R}^{m\times n}$ with $k\leq r\leq\min\{m,n\}$, where $W$ is full rank, $K(N)<\varepsilon\leq\mathcal{O}\left(\frac{\sigma_{k}(W)}{\mathcal{K}(W)}\right)$ and $H\in\mathbb{R}_{+}^{r\times n}$ is separable with $H^{\top}e\leq e$. If $K(Q)\leq\mathcal{O}\left(\mathcal{K}(W)\varepsilon\right)$, then the column of $X$ with the largest $\ell_{2}$ norm, denoted $x$, satisfies
@@
In this section, we improve the bound of Theorem 2; see Theorem 3 below which is very similar to Theorem 2, but improves the error bound, namely from $\mathcal{O}\left(\varepsilon\mathcal{K}^{2}(W)\right)$ to $\mathcal{O}\left(\varepsilon\mathcal{K}(W)\right)$. Note that it does not improve [17, Theorem 2] in terms of the bound on the noise allowed by SPA which was already of the order of $\varepsilon\leq\mathcal{O}\left(\frac{\sigma_{r}(W)}{\mathcal{K}(W)}\right)$.
@@
To prove Theorem 3, we use a similar proof strategy as that of [20, Theorem 1], which is based on the geometry of the problem, as opposed to matrix norm inequalities used in [17], as explained in [20]. However, [20, Theorem 1] still has a factor proportional to $\mathcal{K}^{2}(W)$ in the error bound, and hence its bound is asymptotically similar to that of Theorem 2. Moreover, [20, Theorem 1] claims that it can replace $\sigma_{r}(W)$ in Theorem 1 by $\sigma_{r-1}(W)$ in the bound on the noise and in the error bound. However, this is not possible, as exemplified by the following counter example. Suppose in fact that $W=\begin{pmatrix}0&0\\ 1&0\end{pmatrix}$, $r=2$ and $X=\begin{pmatrix}0&\varepsilon&0\\ 1&1&0\end{pmatrix}$ for a small enough $\varepsilon\geq 0$. Notice that $X=WH+N$ for a separable and column stochastic $H=\begin{pmatrix}1&1&0\\ 0&0&1\end{pmatrix}$ and with $K(N)=\varepsilon$. Then SPA identifies the first two columns of $X$ as the vertices of the simplex, and the error is at least $1$ for any $\varepsilon$, that does not coincide with the estimation $\mathcal{O}(\varepsilon K(W)^{2}/\sigma_{r-1}(W)^{2})=\mathcal{O}(\varepsilon)$.
@@
The consequence of Theorem 3 are the following.
  • The first step of SPA is more robust than the next ones. The reason is that the projection step might lead to an increase of the error bound by a factor of $\mathcal{K}(W)$, so that Theorem 1 becomes tight after the first step of SPA. We will prove this in Section 5 with an example.
@@ 3.2 SPA when $r=2$
$\implies\lambda$ $\leq\frac{2\varepsilon}{\|\widetilde{w}_{2}\|-\|\widetilde{w}\|}\leq\frac{2\varepsilon}{(1-\frac{1}{2}\frac{1}{1-c\gamma(c)})\|\widetilde{w}_{2}\|}$
@@
$=\left(1+\beta(c)\mathcal{K}(W)\right)\varepsilon\leq\left(1+\beta(c)\right)\mathcal{K}(W)\varepsilon$
@@ 4.1 First two steps of T-SPA
$\varepsilon\leq c\sigma_{r}(W)\mathcal{K}(W)^{-1}\leq\frac{4c}{3-2\sqrt{2}}\sigma_{r-1}(\hat{W})\mathcal{K}(\hat{W})^{-1}\leq\sigma_{r-1}(\hat{W})\mathcal{K}(\hat{W})^{-1}/16,$
@@ 5 Tightness of the error bound in Theorem 1 when $r\geq 3$
In this section, we prove that the error bound in Theorem 1 is tight when $r\geq 3$ (up to the hidden constant factors), which was an important open question regarding SPA [14, p. 231].
@@
For $r\geq 3$, the second step of SPA might have an error in $\mathcal{O}\left(\varepsilon\mathcal{K}^{2}(W)\right)$. This implies that the error bound in Theorem 1 is tight for $r\geq 3$, up to constant terms.

### v2

@@ 1 Introduction
Countless algorithms have been developed for separable SSMF; see, e.g., [15, Chap. 7] and the references therein. Among them, a workhorse algorithm is the successive projection algorithm (SPA) [3] which is described in Algorithm 1. Geometrically, when $\operatorname{rk}(W)=r$, $\operatorname{conv}(W)$ is an $(r-1)$-dimensional simplex with $r$ vertices embedded in a higher $m$-dimensional space. The first step of SPA extracts the point with the largest $\ell_{2}$ norm. For any polytope (such as a simplex), such a point is always a vertex. Then SPA projects all data points on the orthogonal complement of that extracted point (so that this point is projected onto zero), reducing the dimension of the simplex by one, and continuing the process $(r-1)$ times.
@@
3:        $p\in\operatorname{argmax}_{j}\|R(:,j)\|$.    % Selection step
@@
5:        $R\leftarrow\left(I-\frac{{R(:,p)}R(:,p)^{\top}}{\|{R(:,p)}\|^{2}}\right)R$.      % Projection step
@@
SPA should not be implemented as in Algorithm 1, this would be particularly ineffective for large sparse matrices, as the residual would become quickly dense (especially if the first extracted columns is dense). It can be implemented in $\mathcal{O}(r\operatorname{nnz}(X))$ operations (essentially, $r$ matrix-vector products), where $\operatorname{nnz}(X)$ is the number of non-zero entries of $X$, using the fact that $\|(I-uu^{\top})y\|^{2}=\|y\|^{2}-(u^{\top}y)^{2}$ where $u$ has unit norm and $y$ is any vector [19]. See https://gitlab.com/ngillis/robustSPA for a MATLAB implementation.
SPA has a long history, as the index set of columns it extracts corresponds to that of QR with column pivoting [8]. However, it was introduced in the context of separable SSMF by Araújo et al. [3], and rediscovered many times under different names; see [15, p. 229] for an historical overview.
@@
[19, Theorem 3] Let $X=WH+N\in\mathbb{R}^{m\times n}$ where $H\in\mathbb{R}^{r\times n}_{+}$ is a separable matrix satisfying $H^{\top}e\leq e$, with $e$ the vector of all one of appropriate dimension, and let
$\varepsilon=\max_{j}\|N(:,j)\|\leq\mathcal{O}\left(\frac{\sigma_{r}(W)}{\sqrt{r}\mathcal{K}^{2}(W)}\right),$ (Bound on the noise)
where $\sigma_{r}(W)$ is the $r$th singular value of $W$, and $\mathcal{K}(W)=\frac{K(W)}{\sigma_{r}(W)}$ with $K(W)=\max_{k}\|W(:,k)\|$ is a measure of the conditioning [^2] Note that $\mathcal{K}(W)$ is smaller than the condition number of $W$, $\frac{\sigma_{1}(W)}{\sigma_{r}(W)}$. In fact, recall that $\sigma_{1}(W)=\|W\|\geq K(W)\geq\sigma_{1}(W)/\sqrt{r}$, where $r$ is the number of columns in $W$. of $W$. Then SPA (Algorithm 1) returns a set of indices $\mathcal{J}$ such that
$\max_{1\leq k\leq r}\min_{j}\|W(:,k)-X(:,\mathcal{J}_{j})\|\leq\mathcal{O}\left(\varepsilon\mathcal{K}^{2}(W)\right).$ (Error bound)
@@
  • Section 5. For $r\geq 3$, we show that the bound of Theorem 1 from [19] with error bound in $\mathcal{O}\left(\varepsilon\mathcal{K}^{2}(W)\right)$ is tight (Theorem 7). We do this by providing a family of data matrices that achieve this worst-case bound already at the second step of SPA. This answers an open question about the tightness of the SPA error bound of Theorem 1 [15, p. 231].
@@
Given a vector $x\in\mathbb{R}^{m}$, we denote $\|x\|$ its $\ell_{2}$ norm. Given a matrix $X\in\mathbb{R}^{m\times n}$, we denote $X^{\top}$ its transpose, $X(:,j)$ its $j$th column, $X(i,:)$ its $i$th row, $X(i,j)$ its entry as position $(i,j)$, $\|X\|=\sigma_{\max}(X)$ its $\ell_{2}$ norm which is equal to its largest singular value, $\|X\|_{F}^{2}=\sum_{i,j}X(i,j)^{2}$ its squared Frobenius norm, $\sigma_{r}(X)$ its $r$th singular value, $\sigma_{\min}(X)=\sigma_{\min(m,n)}(X)$ its smallest singular value, $\operatorname{rk}(X)$ its rank, $K(X)$ $=$ $\max_{j}\{\|X(:,j)\|\}$, and $\mathcal{K}(X)=K(X)/\sigma_{\min}(X)$. For $k>\min\{m,n\}$, by convention, $\sigma_{k}(X)=0$. Note that
@@
and $\sigma_{n}(X)\sqrt{2}\leq\|X(:,i)-X(:,j)\|$ for every $i\neq j$, where $P_{u}(v)$ is the projection of $v$ on the subspace orthogonal to $u$. We denote $e_{k}$ the $k$th unit vector, $e$ the vector of all ones of appropriate dimension, $I_{r}$ the identity matrix of dimension $r$. The set +$\mathbb{R}^{m\times n}_{+}$ denotes the $m$-by-$n$ component-wise nonnegative matrices. A matrix $H\in\mathbb{R}^{r\times n}$ has stochastic columns if $H\geq 0$ and $H^{\top}e=e$. It is separable if $H(:,{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{J}}})=I_{r}$ for some index set $\mathcal{J}$ of size $r$.
@@
In the results, we use the notation $\dots\leq\mathcal{O}(f(\bf x))$ to indicate that a quantity is bounded by $f(\bf x)$ up to an absolute multiplicative constant $C$ that does not depend on any of the parameters or variables.
@@ 2 SPA: what do we know?
    The error bound are rather poor, depending on the squared of the condition number of $W$. Moreover, it is not known whether this bound is tight [15, p. 231].
@@ 2.1 Drawback 1: rank-deficient case
3:        $p\in\operatorname{argmax}_{j}\|R(:,j)\|$.                 % Selection step
@@
8:              $R\leftarrow\left(I-\frac{{R(:,p)}R(:,p)^{\top}}{\|{R(:,p)}\|^{2}}\right)R$.      % Projection step
@@
As for SPA (see Remark 1), T-SPA can be implemented in $\mathcal{O}(r\operatorname{nnz}(X))$ operations (essentially, $r$ matrix-vector products). See https://gitlab.com/ngillis/robustSPA for a MATLAB implementation.
@@
    Note that we will always have $\mathcal{K}(W_{t})\leq 2\mathcal{K}(W)$. In fact, $K(W_{t})\leq 2K(W)$ since $\|w_{i}-w_{j}\|\leq\|w_{i}\|+\|w_{j}\|$, and $\sigma_{r-1}(W_{t})\geq\sigma_{r}(W)$ since $W_{t}$ is a rank-one correction of $W$ (by the interlacing singular value theorem).
@@ 3 Improved bounds for the first step of SPA and when $r=2$
Motivations
From a theoretical standpoint, understanding better the behavior of SPA, a workhorse separable NMF algorithm, is a mathematically interesting problem, even when $r=2,3$. From a practical standpoint, rank-2 NMF and/or SPA with $r=2$ have been used successfully to:
  • segment and analyze medical images [25, 35],
  • extract topics in large corpus of documents [23], and
  • cluster pixels in hyperspectral images [16].
In analytical chemistry, where NMF is referred to as the self-modeling curve resolution (SMCR) problem, using $r=3$ is a standard setting [7, 32].
@@ 3.1 First step of SPA
[19, Theorem 2] Let $W\in\mathbb{R}^{m\times k}$, and $Q\in\mathbb{R}^{m\times(r-k)}$ and $X=[W\,\,Q]H+N\in\mathbb{R}^{m\times n}$ with $k\leq r\leq\min\{m,n\}$, where $W$ is full rank, $K(N)<\varepsilon\leq\mathcal{O}\left(\frac{\sigma_{k}(W)}{\mathcal{K}(W)}\right)$ and $H\in\mathbb{R}_{+}^{r\times n}$ is separable with $H^{\top}e\leq e$. If $K(Q)\leq\mathcal{O}\left(\mathcal{K}(W)\varepsilon\right)$, then the column of $X$ with the largest $\ell_{2}$ norm, denoted $x$, satisfies
@@
In this section, we improve the bound of Theorem 2; see Theorem 3 below which is very similar to Theorem 2, but improves the error bound, namely from $\mathcal{O}\left(\varepsilon\mathcal{K}^{2}(W)\right)$ to $\mathcal{O}\left(\varepsilon\mathcal{K}(W)\right)$. Note that it does not improve [19, Theorem 2] in terms of the bound on the noise allowed by SPA which was already of the order of $\varepsilon\leq\mathcal{O}\left(\frac{\sigma_{r}(W)}{\mathcal{K}(W)}\right)$.
@@
To prove Theorem 3, we use a similar proof strategy as that of [22, Theorem 1], which is based on the geometry of the problem, as opposed to matrix norm inequalities used in [19], as explained in [22]. However, [22, Theorem 1] still has a factor proportional to $\mathcal{K}^{2}(W)$ in the error bound, and hence its bound is asymptotically similar to that of Theorem 2. Moreover, [22, Theorem 1] claims that it can replace $\sigma_{r}(W)$ in Theorem 1 by $\sigma_{r-1}(W)$ in the bound on the noise and in the error bound. However, this is not possible, as exemplified by the following counter example. Suppose in fact that $W=\begin{pmatrix}0&0\\ 1&0\end{pmatrix}$, $r=2$ and $X=\begin{pmatrix}0&\varepsilon&0\\ 1&1&0\end{pmatrix}$ for a small enough $\varepsilon\geq 0$. Notice that $X=WH+N$ for a separable and column stochastic $H=\begin{pmatrix}1&1&0\\ 0&0&1\end{pmatrix}$ and with $K(N)=\varepsilon$. Then SPA identifies the first two columns of $X$ as the vertices of the simplex, and the error is at least $1$ for any $\varepsilon$, that does not coincide with the estimation $\mathcal{O}(\varepsilon K(W)^{2}/\sigma_{r-1}(W)^{2})=\mathcal{O}(\varepsilon)$.
@@
The consequences of Theorem 3 are the following.
  • The first step of SPA is more robust than the next ones. The reason is that the projection step might lead to an increase of the error bound by a factor of $\mathcal{K}(W)$, so that Theorem 1 becomes tight after the first step of SPA. We will prove this in Section 5 with an example. From a practical point, this is an interesting information: we know that the first extracted vertex by SPA is more reliable than the next ones.
@@ 3.2 SPA when $r=2$
$\implies\lambda$ $\leq\frac{2\varepsilon}{\|\widetilde{w}_{2}\|-\|\widetilde{w}\|}\leq\frac{2\varepsilon}{(1-\frac{1}{2}\frac{1}{1-c\gamma(c)})\|\widetilde{w}_{2}\|}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}.}}$
@@
$=\left(1+\beta(c)\mathcal{K}(W)\right)\varepsilon\leq\left(1+\beta(c)\right)\mathcal{K}(W)\varepsilon{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},}}$
@@ 4.1 First two steps of T-SPA
$\varepsilon\leq c\sigma_{r}(W)\mathcal{K}(W)^{-1}\leq\frac{4c}{3-2\sqrt{2}}\sigma_{r-1}(\hat{W})\mathcal{K}(\hat{W})^{-1}\leq\sigma_{r-1}(\hat{W})\mathcal{K}(\hat{W})^{-1}/16{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}.}}$
@@ 5 Tightness of the error bound in Theorem 1 when $r\geq 3$
In this section, we prove that the error bound in Theorem 1 is tight when $r\geq 3$ (up to the hidden constant factors), which was an important open question regarding SPA [15, p. 231].
@@
For $r\geq 3$, the second step of SPA might have an error in ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Omega}}\left(\varepsilon\mathcal{K}^{2}(W)\right)$. In particular, there exist a full rank matrix $W\in\mathbb{R}^{3\times 3}$ and an absolute constant $c>0$ such that, for any $\varepsilon$ in the interval $(0,c)$, there exist $H_{\varepsilon}\in\mathbb{R}^{3\times 4}$ separable and column stochastic, and $N_{\varepsilon}\in\mathbb{R}^{3\times 4}$ with $K(N_{\varepsilon})\leq\varepsilon$, for which the error of the second step of SPA applied to $WH_{\varepsilon}+N_{\varepsilon}$ is larger than $c\varepsilon\mathcal{K}^{2}(W)$. This implies that the error bound in Theorem 1 is tight for $r\geq 3$, up to constant terms.